OpenStudy (ineedhelpnow):

solve (xy-3)^2 using FOIL

OpenStudy (rhystic):

(xy-3)(xy-3) xy^2 - 3xy -3xy +9 xy^2 - 6xy +9 basically you take the first xy multiply it by both parts in the second parenthesis and then do the same with the 3

OpenStudy (anonymous):

(xy - 3)(xy - 3) Foil: First Inside, Ouside Last inside -------- outside xy*xy -3xy -3xy + 9 (x^2)(y^2) - 6xy + 9

OpenStudy (anonymous):

x^2y^2-(2)(3xy)+9

OpenStudy (lgbasallote):

(xy-3)(xy-3) = x^2y^2 - 3xy -3xy +9 = x^2y^2 -6xy + 9

OpenStudy (ineedhelpnow):

Thanks guys!